please help me with this problem.
2/7x2&1/8x3=
The answers are.....6 & 1/28, 9/4, 1 & 23/28, or none of these.

Answer by elima(1433)

2/7x2&1/8x3=
Begin by making the 2 1/8 an improper fraction = 17/8
Then put 3 in fraction form, 3/1
So now we have:
2/7*17/8*3/1=
Now we just multiply all the numerators together, and all denominators.
2*17*3=102
7*8*1=56
So you have: 102/56 and when you simplify you get:
1 46/56 simplify again = 1 23/28
Hope this helps =)